does.google com Mr. Smith is selling roses for valentines day. He sells each dozen for $24.95. So far this month, he has sold 93 dozens. Let x represent the number of additional dozen of roses he expect to sell for the season. Let s(x) represent the amount of money he expect to receive for the entire seasons sales. Write an equation for s(x).* Your answer Solve the system of equations below for x, y, z. * -a+3b-6c 20 a+4b-c= 15 〇@s.2, b 24,0-.1 The materials for each dozen of roses cost a total of $7.75 Write a function S(x) for the profit Mr. Smith earn when he sellx
Answer #1

each dozen is sold for $ 24.95

total dozens = 93

x represents each dozen sold

so, S(x) = 93*24.95 + 24.95 x

S(x) = 24.95x + 2320.35

2)  -a + 3b - 6c = 20

a + 4b - c = 15

-7b - 5c = -23

from equation 3

-7b = -23 + 5c

b = 23/7 - 5/7 c

plugging this value of b into equation 1 and 2

-a + 3( 23/7 - 5/7 c ) - 6c = 20

a + 4(23/7 - 5/7c ) - c = 15

solving we get

a = - 2 , c = -1

plugging a and c value into equation 3

b = 4

hence, solution is

a = - 2 , b = 4 , -1

2nd option is correct